The probability that a person owns a Adele song is .75, that a person owns a Blake Shelton song is .65, and that a person owns both is .23. Find the probabilitiy that a person owns either song but not both.
Answer by ikleyn(52778)

.
The probability that a person owns a Adele song is .75, that a person owns a Blake Shelton song is .65,
and that a person owns both is .23. Find the probability that a person owns either song but not both.
~~~~~~~~~~~~~~~~~~~


The input data represent IMPOSSIBLE situation, which NEVER may happen, since 


    P(A U B) = P(A) + P(B) - P(A n B) = 0.75 + 0.65 - 0.23 = 1.17


is greater than 1.


So the problem is good for only one purpose: to throw it into the GARBAGE BOX.


It is what I recommend you to do immediately.


Have a nice day.